# A kinetic-based regularization method for data science applications

### v1

@@ III.4 Computational Cost
Figure 9: The representative computational cost of the Lagrange multiplier calculations for each training point in the present method is compared with the weight determination step of the Radial Basis Function (RBF) across a range of datasets. It is evident that for asymptotically large $N$ (around $10^{5}$), the present method exhibits better efficiency.
@@
The most computationally intensive step in Algorithm 1 is the calculation of first-order corrections, which involves solving for the Lagrange multipliers (LM). This step requires solving $D$ simultaneous equations at each training point to compute the Lagrange multipliers. In contrast, the RBF method requires solving a set of $N$ coupled algebraic equations (cf. Eq. 4) to compute a set of unknown weights.
@@
In order to establish a fair comparison, we use NumPy’s $linalg$ package to solve the linear equations involved in both the Lagrange multiplier and weight calculations. We perform tests using training sets of increasing sizes, recording computation times for both methods.
@@ Appendix A Calculation of Lagrange Multipliers
We are interested to find an interpolation value at $\bm{x}$. In the expression $\hat{\phi}(\bm{x})=\sum_{i}\phi(\bm{x_{i}})P(\|\tilde{\bm{\xi}}_{i}||,\theta)=\sum_{i}\phi(\bm{x_{i}})P(\|\tilde{\bm{x}}-\bm{x_{i}}\|,\theta)$, we inherently assume a zero correction, i.e., $\tilde{x}=x$. We, however, wish to construct a distribution function that satisfies the first moment condition given in Eq.(7). For a locally linear approximation of a given function $\phi(x)$ in 1D, we have the condition,
$\begin{split}\sum_{i}(x-x_{i})P(\|\tilde{x}^{(k+1)}-x_{i}\|,\theta)=0.\end{split}$ (26)
Here, $k+1$ is the iterative index, the method for which we will discuss below. Simplifying and writing $\tilde{x}=x+\delta x$. For an iterative procedure, it yields $\tilde{x}^{(k+1)}=\tilde{x}^{(k)}+\delta\tilde{x}^{(k)}$, where $\tilde{x}^{(0)}=x$. It can be worked out in a few simple steps that $P$ in Eq. (26) need not be normalized. Hence, we can work here with the non-normalized Gaussian distribution is given by
@@
Simplifying and neglecting $(\delta x)^{2}$ term, and omitting the superscript $(k)$ we obtain a transcendental equation to be solved for $\delta x$ wiz,
@@
Linearizing this equation assuming a small $\delta x$,
@@
$\sum_{i}(x-x_{i})P(\|\tilde{x}-x_{i}\|,\theta)\\ -\delta\tilde{x}\sum_{i}\frac{1}{\theta}(\tilde{x}-x_{i})(x-x_{i})P(\|\tilde{x}-x_{i}\|,\theta)=0.$ (31)
Hence we get a solution,
@@
$\delta\tilde{x}^{(k)}=\theta\frac{\sum_{i}(x-x_{i})P(\|\tilde{x}^{(k)}-x_{i}\|,\theta)}{\sum_{i}\left((\tilde{x}^{(k)}-x_{i})(x-x_{i})P(\|\tilde{x}^{(k)}-x_{i}\|,\theta)\right)}.$ (33)
So,
@@
Once we have the corrected $x$, we use this value for the calculation of distribution function values. In $D$-dimensions where $\bm{x}=x_{\alpha}\,\hat{e}_{\alpha}$ , the same algebra of Eq(30) simplifies to,
$\sum_{i}(x_{\alpha}-x_{\alpha,i})P(\|\tilde{\bm{x}}-\bm{x_{i}}\|,\theta)\\ -\frac{\delta\tilde{x}_{\kappa}}{\theta}\sum_{i}(\tilde{x}_{\kappa}-x_{\kappa,i})(x_{\alpha}-x_{\alpha,i})P(\|\tilde{\bm{x}}-\bm{x_{i}}\|,\theta)=0.$ (35)
Here, $\kappa$ is a repeated index, and the above represents a linear algebra problem.
Let, $P(\|\tilde{\bm{x}}-\bm{x_{i}}\|,\theta)=P_{i}$ for brevity.
In 2D, if we represent a coordinate $\bm{x}_{i}$ as $\bm{x}_{i}=x_{i}\hat{e}_{x}+y_{i}\hat{e}_{y}$, Eq. (10) translates to $\mathcal{A}:\delta\bm{X}=\bm{B}$ problem where,
@@
This can be solved iteratively for a point $\bm{x}$ to obtain the corrections.

### v2

@@ III.4 Computational Cost
Figure 10: The representative computational cost of the Lagrange multiplier calculations for each training point in the present method is compared with the weight determination step of the Radial Basis Function (RBF) across a range of datasets. It is evident that for asymptotically large NN (around $10^{5}$), the present method exhibits better efficiency.
@@
The most computationally intensive step in Algorithm 1 is the calculation of first-order corrections, which involves solving for the Lagrange multipliers (LM). This step requires solving $D$ simultaneous equations at each training point to compute the Lagrange multipliers. In contrast, the RBF method requires solving a set of NN coupled algebraic equations (cf. Eq. 4) to compute a set of unknown weights.
@@
In order to establish a fair comparison, we use NumPy’s linalglinalg package to solve the linear equations involved in both the Lagrange multiplier and weight calculations. We perform tests using training sets of increasing sizes, recording computation times for both methods.
@@ IV Conclusions
Data availability statement
An implementation of the algorithm, alongside scripts and datasets to reproduce the results presented in this work, can be found in the following GitHub repository: https://github.com/agabbana/kinetic-based-regularization
@@ Appendix A Calculation of Lagrange Multipliers
Figure 11: Performance comparison between Gaussian Process Regression (GPR) and the present method in the $6D$ sparse-grid regime, for the Rastrigin function (uncoupled) and Camel function (multiplicatively coupled). Top panel shows the RMSE on a random test set of 10,000 points with varying training set sizes NN. Bottom panel shows the corresponding computation times.
We are interested in finding an interpolated value at a point $\bm{x}$. In the expression $\hat{\phi}(\bm{x})=\sum_{i}\phi(\bm{x_{i}})P(\|\tilde{\bm{\xi}}_{i}||,\theta)=\sum_{i}\phi(\bm{x_{i}})P(\|\tilde{\bm{x}}-\bm{x_{i}}\|,\theta)$, we inherently assume a zero correction, i.e., $\tilde{x}=x$. However, we aim to construct a distribution function that satisfies the first moment condition given in Eq. (7). For a locally linear approximation of a function $\phi(x)$ in 1D, we impose the condition
$\begin{split}\sum_{i}(x-x_{i})P(\|\tilde{x}^{(k+1)}-x_{i}\|,\theta)=0,\end{split}$ (28)
where $k+1$ denotes the iteration index (to be discussed below).
Writing $\tilde{x}=x+\delta x$ and applying an iterative procedure, we have $\tilde{x}^{(k+1)}=\tilde{x}^{(k)}+\delta\tilde{x}^{(k)}$, where $\tilde{x}^{(0)}=x$. It can be shown that $P$ in Eq. (28) does not require a normalization, allowing us to work with the non-normalized Gaussian distribution:
@@
Neglecting terms of order $(\delta x)^{2}$ and omitting the superscript $(k)$, we obtain a transcendental equation for $\delta x$:
@@
Linearizing assuming small $\delta x$,
@@
$\sum_{i}(x-x_{i})P(\|\tilde{x}-x_{i}\|,\theta)\\ -\delta\tilde{x}\sum_{i}\frac{1}{\theta}(\tilde{x}-x_{i})(x-x_{i})P(\|\tilde{x}-x_{i}\|,\theta)=0,$ (33)
which yields the solution
@@
$\delta\tilde{x}^{(k)}=\theta\frac{\sum_{i}(x-x_{i})P(\|\tilde{x}^{(k)}-x_{i}\|,\theta)}{\sum_{i}\left((\tilde{x}^{(k)}-x_{i})(x-x_{i})P(\|\tilde{x}^{(k)}-x_{i}\|,\theta)\right)},$ (35)
and hence
@@
Once we have the corrected $x$, we use this value for the calculation of the values of the distribution function. Extending this to $D$ dimensions where $\bm{x}=x_{\alpha}\hat{e}_{\alpha}$, the algebra in Eq. 32 generalizes to
$\sum_{i}(x_{\alpha}-x_{\alpha,i})P(\|\tilde{\bm{x}}-\bm{x_{i}}\|,\theta)\\ -\frac{\delta\tilde{x}_{\kappa}}{\theta}\sum_{i}(\tilde{x}_{\kappa}-x_{\kappa,i})(x_{\alpha}-x_{\alpha,i})P(\|\tilde{\bm{x}}-\bm{x_{i}}\|,\theta)=0,$ (37)
where, $\kappa$ is a repeated index, and the above represents a linear algebra problem.
Letting $P(\|\tilde{\bm{x}}-\bm{x}_{i}\|,\theta)=P_{i}$ for brevity, in 2D, representing coordinates as $\bm{x}_{i}=x_{i}\hat{e}_{x}+y_{i}\hat{e}_{y}$, Eq. (37) translates to the matrix problem $\mathcal{A}:\delta\bm{X}=\bm{B}$, where
@@
This system can be solved iteratively for each point $\bm{x}$ to obtain the correction $\delta\bm{X}$.
Appendix B Hermite polynomial expansion approach
Figure 12: Optimal temperature estimation for the 2D Franke function using 2000 random sample points ($N=2000$). The left/right panel shows results without/with corrections. Filled markers indicate the minimum values. In both cases, the present method converges in 13 iterations, while MLE requires 46 and 52 iterations, respectively.
Let us consider a target function, $\phi(\bm{x})$, which admits a second-order Hermite expansion Grad (1949):
$\phi(\bm{x})=a_{0}\,\mathcal{H}^{(0)}(\bm{x})+a_{1\,\alpha}\,\mathcal{H}^{(1)}_{\alpha}(\bm{x})+a_{2\,\alpha\beta}\,\mathcal{H}^{(2)}_{\alpha\beta}(\bm{x})+\cdots.$ (39)
Here, $\alpha$ and $\beta$ are representatives of dimensional indices. The coefficients can be computed via projection with respect to a Gaussian weight:
$a_{0}(\bm{x})=\int\exp\left(-\frac{\bm{x}^{2}}{2\theta}\right)\,\phi(\bm{x})\,d\bm{x},$ (40)
$a_{1\,\alpha}(\bm{x})=\int x_{\alpha}\,\exp\left(-\frac{\bm{x}^{2}}{2\theta}\right)\,\phi(\bm{x})\,d\bm{x},$
$a_{2\,\alpha\beta}(\bm{x})=\int(x_{\alpha}\,x_{\beta}-\theta\delta_{\alpha\beta})\exp\left(-\frac{\bm{x}^{2}}{2\theta}\right)\,\phi(\bm{x})\,d\bm{x}.$
Note that as the third-order contribution is ignored, errors will be proportional to the third Hermite coefficient. Then for a given point of interest $\bm{x}_{i}$}, our approximate projection is given by,
$\hat{\phi}(\bm{x}_{i})=\int\phi(\bm{x})P(\bm{\xi}_{i},\theta)\,d\bm{x},$ (41)
[…]
$\hat{\phi}({\bf x_{i}})=\int\big{(}$ $a_{0}\,{\cal H}^{(0)}({\bm{x}})+a_{1\,\alpha}\,{\cal H}^{(1)}_{\alpha}({\bm{x}})$
$+a_{2\,\alpha\beta}\,{\cal H}^{(2)}_{\alpha\beta}({\bm{x}})+\cdots\big{)}P(\bm{\xi}_{i},\theta)\,d\bm{x}$ (42)
Here, the constants are not a function of $\bm{x}$ and can be taken out of the integral. Therefore we have,
$\tilde{\phi}({\bm{x}_{i}})=a_{0}\int{\cal H}^{(0)}({\bm{x}})P(\bm{\xi}_{i},\theta)d\bm{x}\\ +a_{1\,\alpha}\int{\cal H}^{(1)}_{\alpha}({\bm{x}})P(\bm{\xi}_{i},\theta)\,d\bm{x}\\ +a_{2\,\alpha\beta}\int\,{\cal H}^{(2)}_{\alpha\beta}({\bm{x}})P(\bm{\xi}_{i},\theta)\,d\bm{x}+\cdots$ (43)
Notice that the orthogonal Hermite polynomials are ${\cal H}^{(0)}=1$, ${\cal H}^{(1)}=x_{\alpha}$, ${\cal H}^{(2)}=x_{\alpha}x_{\beta}-\theta\delta_{\alpha\beta}$ etc. Substituting these forms in the equation above, we get
$\hat{\phi}({\bm{x}_{i}})=a_{0}\int P(\bm{\xi}_{i},\theta)\,d\bm{x}+a_{1\,\alpha}\int x_{\alpha}P(\bm{\xi}_{i},\theta)\,d\bm{x}\\ +a_{2\,\alpha\beta}\int(x_{\alpha}x_{\beta}-\theta\delta_{\alpha\beta})P(\bm{\xi}_{i},\theta)\,d\bm{x}+\cdots$ (44)
From here, we can see that if the first-order correction, that is, the Lagrange multiplier $\tilde{\bm{\xi}_{i}}$ in place of the naive $\bm{\xi}_{i}$ (or the point of interpolation $\bm{x}_{i}$) is used, we get.
$\phi({\bm{x}_{i}})=a_{0}+a_{1\,\alpha}x_{i\,\alpha}+a_{2\,\alpha\beta}(x_{i\,\alpha}x_{i\,\beta}+\theta\delta_{\alpha\beta})$
$-a_{2\,\alpha\beta}\int\theta\delta_{\alpha\beta}P(\bm{\tilde{\xi}}_{i},\theta)d\bm{x}$
$\implies\phi({\bm{x}_{i}})=a_{0}+a_{1\,\alpha}x_{i\,\alpha}+a_{2\,\alpha\beta}x_{i\,\alpha}x_{i\,\beta}+\cdots$ (45)
Hence, we get the exact Hermite representation up to the first order here. It is easy to see that if the second-order correction, i.e, the self-correction is implemented as well, we retrieve the exact Hermite representation up to second order, namely
$\phi({\bm{x}_{i}})=a_{0}+a_{1\,\alpha}x_{i\,\alpha}+a_{2\,\alpha\beta}\left(\,x_{i\,\alpha}x_{i\,\beta}-\theta\delta_{\alpha\beta}\right)+\cdots$
$\text{or, }\phi({\bm{x}_{i}})=a_{0}\mathcal{H}^{(0)}+a_{1\,\alpha}\mathcal{H}^{(1)}+a_{2\,\alpha\beta}\mathcal{H}^{(2)}+\cdots.$ (46)
Hence we see that the proposed corrections work in the Hermite space as well, recovering the expansion up to second-order at the interpolation points. The present formulation can be interpreted as the discrete limit of the analysis presented above.
Appendix C Accuracy and Performance for Sparse Data
NN
No
Corrections
No
Corrections (MLE)
All
Corrections
All
Corrections (MLE)
$t_{\text{MLE}}/t_{\text{present}}$ $\theta/d^{2}_{\text{typ}}$ $\theta/d^{2}_{\text{typ}}$ (MLE)
100 $4.89\times 10^{-2}$ $4.85\times 10^{-2}$ $38\times 10^{-2} 3.3$$37\times 10^{-2}$ 2.83 0.65 0.73
500 $2.35\times 10^{-2}$ $2.33\times 10^{-2}$ $1.08\times 10^{-3}$ $1.07\times 10^{-2}$ 14.06 0.69 0.76
1000 $1.31\times 10^{-2}$ $1.31\times 10^{-2}$ $4.07\times 10^{-3}$ $4.07\times 10^{-3}$ 9.75 0.73 0.76
2000 $8.10\times 10^{-3}$ $8.07\times 10^{-3}$ $2.13\times 10^{-3}$ $2.14\times 10^{-3}$ 11.05 1.54 1.43
4000 $6.41\times 10^{-3}$ $6.41\times 10^{-3}$ $8.97\times 10^{-4}$ $8.98\times 10^{-4}$ 8.42 1.50 1.52
8000 $4.41\times 10^{-3}$ $4.37\times 10^{-3}$ $81\times 10^{-4} 6.6$$68\times 10^{-4}$ 11.38 3.05 2.24
12000 $3.47\times 10^{-3}$ $3.45\times 10^{-3}$ $4.00\times 10^{-4}$ $3.99\times 10^{-4}$ 22.62 1.15 1.41
Table 3: Comparison of hyperparameter optimization ($\theta_{\rm optimal}$) between the present method and Maximum Likelihood Estimation with Conjugate Gradient (MLE-CG), based on validation RMSE and training time. Results are obtained using randomly sampled data from the 2D Franke function.
To assess performance in the sparse data regime, particularly for higher-dimensional problems, we analyze the case where the training set size is small ($N<\mathcal{O}(10^{3})$). As a benchmark, we compare our approach with the well-established Additive Kernel Gaussian Process Regression (GPR) method Duvenaud et al. (2011), which is particularly effective for high-dimensional uncoupled functions with sparse training data.
We consider the 6D Rastrigin function, a fully separable benchmark function defined as:
$f_{\text{ras}}(\mathbf{x})=\sum_{i=1}^{6}\left[x_{i}^{2}-A\cos(2\pi x_{i})+A\right],$ (47)
where $\mathbf{x}=(x_{1},x_{2},......,x_{6})\in\mathbb{R}^{6}$, with $A=10$.
For GPR we use the built-in Python library gaussian process. The kernel is constructed as a sum of six Gaussian kernels, one for each dimension:
[…]
Here, each index $i$ corresponds to a spatial dimension, and the kernel is trained by optimizing the individual length scales $l_{i}$ for each Gaussian component. The bounds on the length scales are set between $1\%$ of the distance between two neighboring points and 10 times the domain length, ensuring a reasonable training time. Due to the sparsity and randomness of the sampling grids, the choice of training dataset can strongly affect the results. To mitigate this, all results reported in this study are averaged over 10 independent trials, each using different randomly sampled training sets. The average computation time over three trails is shown in the bottom panel of Fig. 11. As shown in the top panel of Fig. 11, the present method achieves convergence on par with the additive GPR for both functions. However, the bottom panel highlights a key advantage of our method: as the number of training points increases, the computational cost of Additive GPR grows significantly, by a factor of 2 to 3, relative to our approach, making our method more scalable for denser grids. We remark that GPR itself is an area of active research, with recent studies Manzhos and Ihara (2023) showing that it is often possible to bypass the need of invoking MLE to determine the optimal kernel parameters, leading to potential additional speedups.
Appendix D Comparing methods for determining the optimal temperature
In this appendix section we compare the method for determining the optimal value for $\theta$ described in Sec. II.2 against a standard approach such as the Maximum Likelihood Estimate (MLE).
We employ MLE to minimize the validation RMSE which is a function of temperature. In order to ensure the positivity of $\theta$, we consider the following formulation:
$\frac{\partial(RMSE)}{\partial z}=\frac{\partial}{\partial z}\,\left(\frac{\;\Bigl{\langle}\bigl{(}\hat{\phi}(\bm{x})-\phi(\bm{x})\bigr{)}^{2}\Bigr{\rangle}^{\tfrac{1}{2}}}{\max|\phi(\bm{x})|}\right),$ (49)
where $z=\log(\theta)$. We solve the minimization problem using the Conjugate Gradient (CG) method.
We assess both methods in terms of training time and final validation accuracy. For a fair comparison, the CG optimizer is initialized using the same initial guess from Eq. 16.
Figure 12 shows the error landscape as a function of $\theta$, showing that both methods, whether including corrections or not, converge to similar minima. However, in this example the MLE-CG optimization requires more iterations.
Detailed training results are summarized in Table 3. To quantify computational cost, we define the time ratio
$\tau=\frac{t_{\rm MLE}}{t_{\rm present}},$ (50)
which compares the time taken by MLE to that of the present method (with both first- and second-order corrections). We observe that MLE can be up to an order of magnitude more expensive.
The values of $\theta/d^{2}_{\rm typ}$ in the table indicate the effective kernel width relative to the local scale set by the $k$-nearest neighbor distance ($k=5$). It is worth remarking that the proposed method builds a list of candidate $\theta$ values based solely on domain coverage (see Eq. 18), independent of validation error. Despite this, the validation RMSE achieved is comparable to that of MLE across all tested datasets.
NN Without Correction With Correction
Present MLE Present MLE
100 $6.16\times 10^{-2}$ $6.38\times 10^{-2}$ $4.97\times 10^{-2}$ $5.08\times 10^{-2}$
500 $2.21\times 10^{-2}$ $2.19\times 10^{-2}$ $9.06\times 10^{-3}$ $8.99\times 10^{-3}$
1000 $1.46\times 10^{-2}$ $1.47\times 10^{-2}$ $4.25\times 10^{-3}$ $4.23\times 10^{-3}$
2000 $9.11\times 10^{-3}$ $9.06\times 10^{-3}$ $2.69\times 10^{-3}$ $2.68\times 10^{-3}$
4000 $6.54\times 10^{-3}$ $6.54\times 10^{-3}$ $1.28\times 10^{-3}$ $1.28\times 10^{-3}$
8000 $44\times 10^{-3} 4.4$$44\times 10^{-3}$ $1.00\times 10^{-3}$ $9.90\times 10^{-4}$
12000 $3.57\times 10^{-3}$ $3.59\times 10^{-3}$ $7.91\times 10^{-4}$ $7.55\times 10^{-4}$
Table 4: Comparison of RMSE on deployment data between the present method and MLE-CG for randomly sampled 2D Franke function. Values are reported both with and without corrections.
For completeness, both methods are evaluated on an independent test set of $10000$ uniformly spread points. Predictions obtained using the respective optimal temperatures (from both methods) are reported in Table 4. The test performance confirms that, even without relying on validation error, the present method yields nearly equivalent accuracy at a fraction of the computational cost.